OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts on the graph of the equation. y=x^2+6x-1

OpenStudy (anonymous):

do you know the formula?

OpenStudy (anonymous):

no. unless its y=mx+b=c

OpenStudy (anonymous):

no, the quadratic formula is...\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

okay so it's \[x= -6\pm \sqrt{6^2-4(10(-1)}/ 2(1)\]

OpenStudy (anonymous):

sorry no 10 in the equation.

OpenStudy (anonymous):

@HawkCrimson

OpenStudy (anonymous):

the 10 should be a 1 but yeah thats right

OpenStudy (anonymous):

then it's -6+- square root of 36- (-4)/ 2

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

then -6+- sqrt 40 then how do you go from there?

OpenStudy (anonymous):

dont forget that its over 2 still.

OpenStudy (anonymous):

\[\frac{ -6\pm \sqrt{40} }{ 2 }\]

OpenStudy (anonymous):

so you would work it out two ways...you would work it out adding the sqrt40 and subtracting the sqrt40

OpenStudy (anonymous):

yes thank you for reminding me.

OpenStudy (anonymous):

how would you subtract or add 40?

OpenStudy (anonymous):

you would either find the actual sqrt or simplify it.

OpenStudy (anonymous):



OpenStudy (anonymous):

so that leaves us 6+- 2squareroot of 2/ 2

OpenStudy (anonymous):